Direct Labor Cost Budget Ace Racket Company manufactures two types of tennis rackets, the Junior and Pro Striker model...

  1. Direct Labor Cost Budget

    Ace Racket Company manufactures two types of tennis rackets, the Junior and Pro Striker models. The production budget for July for the two rackets is as follows:

    Junior Pro Striker
    Production budget 8,900 units 21,600 units

    Both rackets are produced in two departments, Forming and Assembly. The direct labor hours required for each racket are estimated as follows:

    Forming Department Assembly Department
    Junior 0.25 hour per unit 0.5 hour per unit
    Pro Striker 0.35 hour per unit 0.65 hour per unit

    The direct labor rate for each department is as follows:

    Forming Department $17 per hour
    Assembly Department $10 per hour

    Prepare the direct labor cost budget for July.

    Ace Racket Company
    Direct Labor Cost Budget
    For the Month Ending July 31
    Forming Department Assembly Department
    Hours required for production:
    Junior
    Pro Striker
    Total
    Hourly rate x$ x$
    Total direct labor cost $ $
Answer #1
Forming Assembly
Hours required
Junior (8900*0.25) = 2225 (8900*0.5) = 4450
Pro striker (21,600*0.35) = 7560 (21,600*0.65) = 14040
Total 9785 18490
Hourly rate 17 10
Total 166,345 184,900